# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ IV.1 Preliminaries
Note that this statement is graph independent, and is only an average; individual trajectories will of course display substantial variations. It is a rederivation of a familiar result for dense hypergraphs with Gaussian distributed constraint energies Derrida (1980); Baldwin and Laumann (2018), but is applicable in much broader contexts and is easy to confirm numerically. Throughout this work, we rescale all problems by a multiplicative constant so that the ground state energy is $-N$.
IV.2 Paramagnet to spin glass transition scaling for MAX-3-XORSAT
Before proceeding to our main calculation, we note that, from this result, we can predict the typical case difficulty scaling of finding the ground state with AQC or QAOA, for random hypergraphs as defined in section II.4. To predict the phase transition rate, we will use high order perturbation theory to compute the perturbative dressings to the problem ground state; call this state $\left|G_{D}\right>$. Using fourth order perturbation theory, for uniform transverse field strength $\kappa$, the energy of the dressed ground state is, again on average [^4] In this step the graph structure of $H_{P}$ can potentially be important. This is because the energy denominator $2p$, the average cost per flip just a few flips away from $G$, is more sensitive to the details of the graph than the energy cost many flips away, which can in turn shift $\kappa_{c}$. We thus do not claim this result is valid for all graphs. However, we find that for our random PPSPs the transition is consistently near the $\kappa_{c}$ we predict here and the scaling of the uniform field result closely matches our prediction.
$E_{\mathrm{GS}}\simeq-N\left(1+\frac{\kappa^{2}}{2p}+\frac{\kappa^{4}}{8p^{3}}+...\right)\,.$ (12)
This crosses the energy of the paramagnetic state, which is $-N\kappa$ and minimally perturbed by the problem Hamiltonian, at $\kappa_{c}\simeq 1.29$ for our family of $p=3$ PPSPs. The splitting $2\Omega_{0}$ at the phase transition is expected to be proportional to the overlap of the uniform superposition state $\left|S\right>$, the ground state of $H_{D}$, with the dressed state $\left|G_{D}\right>$:
$\Omega_{0}\propto\left<S|G_{D}\right>.$ (13)
@@ IV.2 Paramagnet to spin glass transition scaling for MAX-3-XORSAT
To compute $\Omega_{0}$, we must thus compute the perturbative corrections to $\left|G_{D}\right>$, to high orders. Again using Eq. (11), if we let $\left|G^{(i,j,k...)}\right>\equiv X_{i}X_{j}X_{k}...\left|G\right>$, and $\tilde{E}_{\mathrm{avg}}\left(k\right)\equiv E_{\mathrm{avg}}\left(k\right)-E_{G}$, at the transition point we have

### v3

@@ IV.1 Cost per flip curve
Note that this statement is hypergraph independent, and is only an average; individual trajectories will of course display substantial variations. It is a rederivation of a familiar result for dense hypergraphs with Gaussian distributed constraint energies Derrida (1980); Baldwin and Laumann (2018), but is applicable in much broader contexts and is easy to confirm numerically.
We do caution however that Eq. 18 is only asymptotically exact, owning to the discretization into single flips in Eq. 17. At small $N$, deviations from it can be meaningful, particularly for larger $p$; see appendix A for more details. Throughout this work, we rescale all problems by a multiplicative constant so that the ground state energy is $-N$. This rescaling assumes that we can correctly guess the fraction $\epsilon$ of unsatisfied constraints in $G$, as mentioned earlier.
Figure 4: Visualization of the composed spectral filters in Eq. 12, as a function of $m$ random flips away from the ground state $G$, normalized so that the ground state energy is $H_{\mathrm{cost}}(m=0)/N=-1$. Blue curve: unfiltered, hypergraph-independent mean energy Eq. 18. Gold curve: same energy with a nonlinear fold (Eq. 8) applied, with $A=1$ and $x=2$. Green curve: applies a spectral warping filter (Eq. 9) after the nonlinear fold, with $w=0.6$; it is this cost function which is optimized in the first stage of our novel multi-stage optimization algorithm.
IV.2 Multi-stage filtered optimization performance analysis
IV.2.1 First stage scaling
With this result in hand, we can now predict the scaling of MSFO, beginning with the first stage. To demonstrate that we find the ground state of $H_{D}+H_{\mathrm{cost}}^{x,w}$, we must predict the scaling of the quantum phase transition that occurs during the ramp up of the cost function Hamiltonian. Such transitions are governed by energetics of flip sequences in the neighborhood of $G$. The cost per flip a few flips away from the ground state determines the location of the transition (as a function of $\kappa$ and $t/T_{1}$). For example, in the unfiltered problem,
$E_{G,D}\simeq-N\left(1+\frac{\kappa^{2}}{2p}+\frac{\kappa^{4}}{8p^{3}}+...\right)\,.$ (19)
This crosses the energy of the paramagnetic state, which is $-N\kappa$ and minimally perturbed by the problem Hamiltonian, at $\kappa_{c}\simeq 1.29$ for our family of $p=3$ PPSPs. In contrast, the high energy structure of states far from $G$ can significantly influence the scaling of the minimum gap, as it is the overlap of the two competing states that determines the width of the resulting avoided crossing.
We first consider the low energy properties of $H_{\mathrm{cost}}^{x,w}$ in Eq. 12, as the precise location of the transition for any given $H_{P}$ will determine the magnitudes of the high-energy wavefunction corrections that ultimately set whether or not the system can successfully make the transition into the ground state of $H_{D}+H_{\mathrm{cost}}^{x,w}$ in time $T_{1}$. As stated in our core conjecture, we are concerned with approximating random instances from the space of problems where classical approximation is hard both formally and, we presume, in practice: almost-satisfiable instances ($\epsilon\ll 1$) with high densities of constraints $d_{C}\gg 1$. We can therefore assume a random hypergraph, and if we draw constraints from the set of $\binom{N}{3}$ unique triples (for a total of $d_{C}N$ constraints), then the distribution of the number of constraints each variable participates in is well approximated as Poissonian, with mean $3d_{C}$ constraints per variable. Since the hardest instances to approximate are those where $d_{C}$ is large but most constraints are satisfied in $G$, we can therefore infer that the local distribution of the cost per flip away from $G$ is also (approximately) Poissonian, [^5] Formally the distribution is not exactly Poissonian, owing to the fact that no triples are repeated in the hypergraph. At large $d_{C}$ we expect this will further suppress the tail of the Poisson distribution, as if we construct the hypergraph by drawing constraints randomly one by one, the number of unique additional constraints to draw decreases with local degree. If anything this will further reduce the local degree variance, and it is the suppression of local variance with $d_{C}$ which is important to our claims. with mean 6 (reflecting the normalization of ground state energy $-N$) and variance $2/\left(d_{C}\left(1-2\epsilon\right)\right)$. Large (e.g. extensive in $N$) fractional fluctuations in the cost per flip, are, in other words, highly suppressed for random approximation-hard instances. Obvious potential exceptions to this are instances with either small $d_{C}$ or larger $\epsilon$, but both of those conditions push us away from the realm of approximation hard instances, at least in practice.
That said, our goal is to confidently predict the structure of the phase transition, and even small (percentage-wise) variations in the ratio of the coefficients of $H_{D}$ and $H_{\mathrm{cost}}^{x,w}$ can meaningfully impact the scaling of the transition. The structure of the composed spectral filters Eq. 12 is chosen to further wash out the influence of the details of $H_{P}$ on the transition into the ground state of $H_{D}+H_{\mathrm{cost}}^{x,w}$. Specifically, the when $x=2$ the mean classical cost per flip cost drops to $O\left(1/N\right)$ about $G$, and $O\left(1\right)$ energies are not reached until $O\left(\sqrt{N}\right)$ classical flips, at which point superpolynomially many states are already contributing to the many-body wavefunction. One can also, for instance, choose $x=3$ in the filter function Eq. 8, at which point the mean classical excitation energy near $G$ is $O\left(1/N^{2}\right)$ and $O\left(1\right)$ energies are not reached until $O\left(N^{2/3}\right)$ flips; the performance of these two formulations is qualitatively similar at reasonably large $d_{C}$. Given a filtered local excitation energy which decays as $O\left(1/N\right)$ or faster, and the exponential suppression of extensive variations in the local excitation energy for random $H_{P}$ drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$, we can assert the following:
Claim 1 (regularity of phase transitions under spectral filtering): for a given choice of $x\geq 2$ and $w<1$ in Eqs. 8-12, and a unit coefficient in front of $H_{\mathrm{cost}}^{x,w}$ so that the energy difference between $G$ and random configurations is $N$, let $\kappa_{c}$ be the average transverse field strength where the phase transition from the paramagnet $\left|S_{D}\right>$ into the dressed problem ground state $\left|G_{D}\right>$ occurs, as determined by the position of the minimum gap. Then, with very high probability, but not guaranteed for all instances, for a given random instance $H_{P}$ drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$, the transverse field strength $\kappa$ where the phase transition occurs will be equal to $\kappa_{c}$ up to corrections that scale as $O\left(1/N^{d}\right)$ for some $d$ that depends on $x$ and $w$.
To further motivate this claim, we make the following observation. The critical transverse field strength $\kappa_{c}$ is defined to be the value of $\kappa$, for unit prefactor $H_{\mathrm{cost}}^{x,w}$, where $E_{S,D}=E_{G,D}$. Since $S\right|H_{\mathrm{cost}}\left|S\right>=0$ and $G\right|H_{D}\left|G\right>=0$, the shifts to these two energies that set $\kappa_{c}$ come entirely from the higher order corrections to each state induced by the other Hamiltonian. And if we use second order perturbation theory as an intuitive guide (but not, we emphasize, as a rigorous calculational tool here), we can see that $E_{S,D}$ is very modestly perturbed by $H_{\mathrm{cost}}$, whereas the shifts to $E_{G,D}$ from $H_{D}$ are significant. For simplicity, let $w=1$, and first let $x=1$ as well, so that our filtering has no physical consequence. Respecting our normalization choice that the energy difference between $E_{G}$ and random configurations is $-N$, the value of each $V_{ijk}$ in $H_{P}$ is ±$\pm\left(d_{C}\left(1-2\epsilon\right)\right)^{-1}$, which is $\ll 1$ for the approximation-hard problems we are concerned with. With this fact in hand, we can readily compute the shift to $E_{S,D}$ using second order perturbation theory, using $H_{D}$ as the base Hamiltonian and $H_{\mathrm{cost}}$ as the perturbation:
$E_{S,D}\simeq-N\left(\kappa+\frac{1}{6\kappa d_{C}\left(1-2\epsilon\right)^{2}}\right).$ (20)
We observe that the correction here is hypergraph independent to this order (for given $\epsilon$ and $d_{C}$) and vanishes as $d_{C}$ becomes large. If we now let $x=2$ (but keep $w=1$), $H_{\mathrm{cost}}$ can be expressed in terms of powers of $H_{P}$, and now contains $O\left(d_{C}^{2}N^{2}\right)$ terms with mean coefficient $N^{-1}\left(d_{C}\left(1-2\epsilon\right)\right)^{-2}$, and the resulting corrections to $E_{S,D}$ scale as $d_{C}^{-2}$ and thus vanish even faster for large $d_{C}$. For non-integer $x$ or $w\neq 1$ the analysis becomes more technical, but the same qualitative result holds: corrections to the energy of the paramagnetic state are very weak in the approximation hard regime. [^6] At small $d_{C}$ or larger $\epsilon$ this may not be the case, of course. For instance, for satisfiable ($\epsilon=0$) three-regular instances, where every spin participates in exactly three constraints, it is easy to show from this analysis that $\kappa_{c}=1$ as the magnitudes of the shifts to the two states are identical.
In contrast, the corrections to $E_{G,D}$ from the transverse field $H_{D}$ are more significant. For $x=w=1$ we find
$E_{G,D}\simeq-N-\kappa^{2}\sum_{j=0}^{N-1}\frac{1}{\Delta E_{1,j}},$ (21)
where $\Delta E_{1,j}$ is the cost to flip spin $j$ starting from $G$ and has mean value $+6$ in our normalization. These corrections do not vanish at large $d_{C}$ and can be quite sensitive to problem structure, at least before any filtering is applied. It is this observation that motivates our filtering choice of $x\geq 2$. While a simple second order perturbative analysis no longer applies since $\Delta E_{1,j}\propto 1/N$ in that limit, making the local cost per flip vanish strongly suppresses the per-instance variation of the classical energy landscape around $G$, and therefore variations in $\kappa_{c}$. Combined with the statistical suppression of variations in the local density of constraints for random instances, we can therefore conclude that per-instance variations in $\kappa_{c}$ will be negligible with high probability, again for random problems drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$ with $d_{C}$ large and $\epsilon$ very small.
We now turn to the high energy structure. As mentioned above, the minimum gap at the phase transition is $\Omega_{0}\propto\left<S_{D}|G_{D}\right>$, where $\left|G_{D}\right>$ is the dressed ground state of the filtered $H_{\mathrm{cost}}$, and $\left|S\right>$ is the uniform superposition ground state [^7] For simplicity, we ignore the perturbative dressings of $\left|S\right>$ by $H_{\mathrm{cost}}^{x,w}$, which may be difficult to calculate in practice, particularly when $x>1$ and $w<1$; these expressions are intended to be illustrative and not exact. The actual prediction of $\Omega_{0}$ for MSFO later in this work relies on a series of well-justified mean field assumptions and does not involve evaluation Eq. 22 directly. of $H_{D}$. Since $H_{D}$ is stoquastic all components of $\left|G_{D}\right>$ will have the same sign, and add constructively in computing $\Omega_{0}$:
$\Omega_{0}\propto\frac{1}{2^{N/2}}\sum_{n=0}^{2^{N}-1}\left<n|G_{D}\right>.$ (22)
As observed in Atia and Aharonov (2019), while the location of the transition is more sensitive to the low energy structure of $H_{\mathrm{cost}}$, its scaling with $N$ is more sensitive to the high energy structure, since most components of $\left|S\right>$ are near $N/2$ flips away from $G$ and thus at energies that are $O\left(N\right)$ above $G$. And while the individual amplitude of any such state in $\left|G_{D}\right>$ will be exponentially small, there are exponentially many of them and they all add constructively in Eq. 22 since they are positive definite for a stoquastic $H_{D}$ Bravyi et al. (2006). This becomes most obvious if we rewrite Eq. 22 by grouping the components of $\left|G_{D}\right>$ by their Hamming distance from $G$ itself:
$\Omega_{0}\propto 2^{-N/2}\sum_{m=0}^{N}\binom{N}{m}\tilde{u}\left(\kappa_{c},x,w,m\right),$ (23)
@@ IV.2.1 First stage scaling
where $\tilde{u}\left(\kappa_{c},x,w,m\right)$ is the average of $n|G_{D}\right>$ over all states $\left|n\right>$ which are $m$ flips away from $G$. For the unfiltered case $x=w=1$ we can do this calculation (at least on average), by computing the perturbative corrections to $\left|G_{D}\right>$, to high orders. Again using Eq. (18), if we let $\left|G^{(i,j,k...)}\right>\equiv X_{i}X_{j}X_{k}...\left|G\right>$, and $\tilde{E}_{\mathrm{avg}}\left(k\right)\equiv E_{\mathrm{avg}}\left(k\right)-E_{G}$, at the transition point we have